Required information Problem 8-3A Asset cost allocation; straight-line depreciation LO C1, P1 (The following information appl3. Using the straight-line method, prepare the December 31 adjusting entries to record depreciation for the 12 months of 2018Journal entry worksheet < 1 2 3 4 Record the year-end adjusting entry for the depreciation expense of Building Note: Enter deJournal entry worksheet < 1 2 Record the year-end adjusting entry for the depreciation expense of Land Improvements 1. Note:Journal entry worksheet < 1 2 3 Record the year-end adjusting entry for the depreciation expense of Land Improvements 2. Note
Answer #1
3
Date General Journal Debit Credit
Dec 31 Depreciation expense—Building 2 27,450 =(624000-75000)/20
Accumulated depreciation—Building 2 27,450
Dec 31 Depreciation expense—Building 3 74,400 =(2262000-402000)/25
Accumulated depreciation—Building 3 74,400
Dec 31 Depreciation expense—Land improvements 1 26,000 =364000/14
Accumulated depreciation—Land improvements 1 26,000
Dec 31 Depreciation expense—Land improvements 2 8,900 =178000/20
Accumulated depreciation—Land improvements 2 8,900
Workings:
Allocation of Purchase Price Appraised Value Percent of
Total Appraised Value
x Total Cost of
Acquisition
= Apportioned Cost
Land 1,860,000 62% x 2,600,000 = 1,612,000
Building 2 720,000 24% x 2,600,000 = 624,000
Land Improvements 1 420,000 14% x 2,600,000 = 364,000
Totals 3,000,000 100% 2,600,000
Land Building 2 Building 3 Land Improvements 1 Land Improvements 2
Purchase Price 1,612,000 624,000 0 364,000 0
Demolition 341,400 0 0 0 0
Land grading 187,400 0 0 0 0
New building (Construction cost) 0 0 2,262,000 0 0
New improvements cost 0 0 0 0 178,000
Totals 2140800 624000 2262000 364000 178000